RC Update and Run plan for Sep 10,2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM

To join the Meeting:  https://bluejeans.com/861439920
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))

Thursday 10 Sept 2020.

General updates

(See d2n Run Plan 2020 for broader status/overview.)

  • Production running on Transverse target; Kin-C, Kin-Z continues.
    • General plan when this is done is to complete Kin-A, Kin-X (partially done already); then move on to Kin-B/Y and/or 1-pass running as time allows.
  • Target Tommy polarization @ 270 is lower than we would like
    • Maximum seen 37.5% [1]; Recent max around 36% (just before tests and Moller this morning) [2]
    • QWP optimization restarted Wed swing; ongoing
    • For comparison, maximum seen with target at 90 deg: 44% (9/5/2020 22:56:38)
  • Extremely trippy beam last night, up to 35--40 trips/hour
    • MCC took beam away at 9am to try to address BLM issues in the North, see if they can recover some cavities, and address issues with the 5th pass separator. Beam back to Hall C around 1:30pm
    • Recovered some RF, we'll see how much it helps [3]
  • Opportunistic Target Studies
    • AFP loss study completed [4]
    • Temperature study @ 270deg completed [5]
  • Moller began at 1:30pm; bailed on it at 3:30pm [6]. We'll try again on Friday at 9:30am.

Notes for the Future

  • Moller measurement for Friday DAY shift (2nd attempt)
  • Beam Studies planned for Tuesday, Sept 15
    • Simona wants access at this time for SOLID cherenkov related work
    • Target work (install spectrum analyzer collimator?)

Pending Issues

  • SHMS s2Y(+/-) scalers stopped counting as of run 11247 (2020-09-03 21:00; last Thursday) [7]
    • ADC/TDC data seem OK, likely a bad module. Will troubleshoot more Friday during Moller.
    • Added scaler check to shift checklist [8]
  • Need spectral analyzer collimator from machine shop
    • Still not ready, Shop says Monday -- that getting pretty late...
    • Junhao will send Brad the shop drawings and we will try to get this done earlier (Hall C Techs?)
  • Target polarization for Tommy @ 270 is low (mid to high 30s)
    • Spectrum analyzer collimator was broken at some point and present optics setup may be allowing transient light to interfere with spectrum measurement and QWP scans
    • JP wonders if the field direction is off for the transverse settings (Bill, Murchhana: see note in Target group chat from late Thursday)
  • Beam energy drifts
    • Hall C 5-pass polarization goes down as energy (HALLC:p) goes up [9]
  • Trip rates
    • Hall A Limited to 140 uA to Reduce ILM0L02 Trips [10]
  • SHMS dipole LN2 loop oscillating [11]
  • pNMR
    • Software / fitting issues? [12]
    • Summary for Tommy at 270 [13], at 90 [14]

Opportunistic Jobs/Access

  • SOLID Cherenkov Test -- Access request [Simona]
    • She would prefer to have access as soon as possible starting this Friday (Sep 11, Day Shift), to change the gas in the Cherenkov Tank during an opportunistic controlled access to the Hall during unexpected down times. NOTE: She may not need tech support, if she can use the ladder to reach fill the tank. \
    • Need ~1 hr without tech support. With tech support, may be 1-2 hrs.
    • Preferably, this work can be done during Day Shift, but it can also be done on Swing/Owl.)

Current Run Plan

Wednesday DAY → Thursday OWL

  • Continue production data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, Transverse/270 polarization with NMR every 5 hrs.

Thursday DAY → Thursday SWING

  • Moller measurement planned for Thursday DAY shift to begin around 9:30am
  • Before handing over to Moller, the target experts will
    • Take NMR/pNMR
    • Perform one of the opportunistic Target Study tests (AFP loss test [B4]?)
  • Wait for Moller to complete and get production beam re-established (late-DAY, early-SWING)
  • Continue production data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, Transverse/270 polarization with NMR every 5 hrs.

Standing Orders

  • If beam goes away for more than 3h, or MCC does operations on the beam that may change the beam tune, then
    • move target to No Target position,
    • verify Harps scans at 3H07A and 3H07B are within spec, and
    • ensure beam is stable (with all locks ON) at 0, 0mm before continuing.
    • Beam recovery procedure has more detailed procedure if needed.
  • New beam positions have been verified and established to be (0,0) mm for all targets.
    • Monitor that beam is at (0,0) using the stripchart on the Wall screen and verify no scraping on the glass cells using the Analysis monitor.
  • Change the IHWP at least every 8 hours when running on glass cells. Call Halls A and B in advance. If there are any issues, call the RC
  • Ensure that all information in the CODA run-start GUIs is correct for every run!
  • Check the position locks are on.
  • Check that CODA is running normally (had one strange hangup lately)
  • Check online replays shortly after starting runs, & full replay results too.
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Follow the updated guidelines in the How-to#Spectrometer_Magnets Spectrometer Magnets How to
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instructions are here: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Useful Links

PolHe3 Polarization Measurements

Misc Links